OpenStudy (anonymous):

what acceleration will be produced when 4000 Newtons of force are applied to a 2000 kilogram car? a. 2 m/s^2 b. 20 m/s^2 c. 2000 m/s^2 d. 8,000,000 m/s^2

OpenStudy (anonymous):

well according to the formula F=ma where F is the force m is the mass and a is the acceleration F/m= a putting the values of F and m 4000/2000=a 2= a thus acceleration 2ms^_2

OpenStudy (anonymous):

F=ma ......so a = F/m = 4000 N/2000 Kg = 2 N/Kg = 2 m/s^2

OpenStudy (anonymous):

Please ask this in the physics group.